OpenStudy (kaylala):

General Identities (in Trigonometry) ADDITION AND SUBTRACTION IDENTITIES FIND THE VALUE OF: sin 15 degrees

OpenStudy (kaylala):

@chmvijay

OpenStudy (anonymous):

\[\sin 15=\sin \left( 45-30 \right)=\sin 45\cos 30-\cos 45\sin 30\] \[\sin 45=\frac{ 1 }{ \sqrt{2} },\cos 45=\frac{ 1 }{ \sqrt{2}},\sin 30=\frac{ 1 }{ 2 },\cos 30=\frac{ \sqrt{3} }{ 2 }\] solve.

OpenStudy (kaylala):

where and how did you get sin (45-30) ???? @surjithayer

OpenStudy (anonymous):

simple math 45-30=15 you can also take 60-45=15

OpenStudy (kaylala):

oh does it give the same answer?

OpenStudy (kaylala):

if i use either? @surjithayer

OpenStudy (anonymous):

yes remember \[\sin 60=\frac{ \sqrt{3} }{ 2 },\cos 60=\frac{ 1 }{ 2 },\]

OpenStudy (kaylala):

how do you know if the ooperation is + or - @surjithayer

OpenStudy (kaylala):

how do you know if the ooperation is + or - at sin (45-30) ??? @surjithayer

OpenStudy (kaylala):

@Preetha @thomaster @ParthKohli

Parth (parthkohli):

Hmm! Are you aware that 15 = 45 - 30?

OpenStudy (anonymous):

we know t-ratios of 0,30,45,60,90

OpenStudy (anonymous):

 divide each by 4 and take the square root.

OpenStudy (kaylala):

@ParthKohli yes i am

OpenStudy (kaylala):

so sin30 becomes\[\frac{ \sqrt{4} }{ 4 }\] ? is it like that? then wht's the point? please help me. i really bad at this and anything related to numbers @surjithayer

OpenStudy (anonymous):

\[\sqrt{\frac{ 4 }{4 }}=\sqrt{1}=1\]

OpenStudy (anonymous):

\[\sqrt{\frac{ 3 }{ 4 }}=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

\[\sin 30=\sqrt{\frac{ 1 }{4 }}=\frac{ 1 }{2 }\]

OpenStudy (kaylala):

i see what about tan?

OpenStudy (anonymous):

\[\sin 90=\sqrt{\frac{ 4 }{4 }}=\sqrt{1}=1\]

OpenStudy (anonymous):

\[\tan 0=\frac{ \sin 0 }{ \cos 0 }=\frac{ 0 }{ 1 }=0\]

OpenStudy (anonymous):

\[\tan \theta =\frac{ \sin \theta }{ \cos \theta } \]

OpenStudy (kaylala):

okay thanks i'll try and figure this out..